HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

CHIRPS: ADIABADIC RAPID PASSAGE USING THZ AND MICROWAVE SOURCES IN MERGED BEAM EXPERIMENTS

\underline{S. HERBERS} , Institute for Molecules and Materials, Radboud University, 6525 AJ Nijmegen, Netherlands, Current address: Univ Paris Est Creteil and Université Paris Cité, CNRS, LISA, F-94010 Créteil, France; S. KUIJPERS, P. SUSTAR, S.Y.T. van de MEERAKKER, Institute for Molecules and Materials, Radboud

University, 6525 AJ Nijmegen, Netherlands

To study bimolecular collisions at low collision energies, we merged two molecular beams of deuterated ammonia. One beam (the primary beam) is controlled using a Stark decelerator, which is a device that exploits the Stark effect to control the velocity of a molecular beam while at the same time acting as a state separator. The second beam passed through a curved hexapole. Both state separators produced nearly pure beams of para − ND_{3} in the rotational ground state J_{|K|}^{±} = 1_{1}^{−} with the sign in the exponent denoting the parity. However, deflection of the primary beam by electric field gradients from the curved hexapole posed a fundamental limitation regarding the phase-space overlap of the merging beams. I will show different attempts to overcome these deflection problems by changing the quantum state of the molecules in the primary beam using microwave (1.6 GHz) and THz (615 GHz) radiation - employing the principles of adiabatic rapid passage - but also by redesigning the experimental setup.

p-number: p252  Submitted on Tue, 27 Jun 2023 15:53:49 +0200